# Stochastic Polyak Step-sizes and Momentum: Convergence Guarantees and Practical Performance

### v1

@@ Abstract
Stochastic gradient descent with momentum, also known as Stochastic Heavy Ball method (SHB), is one of the most popular algorithms for solving large-scale stochastic optimization problems in various machine learning tasks. In practical scenarios, tuning the step-size and momentum parameters of the method is a prohibitively expensive and time-consuming process. In this work, inspired by the recent advantages of stochastic Polyak step-size in the performance of stochastic gradient descent (SGD), we propose and explore new Polyak-type variants suitable for the update rule of the SHB method. In particular, using the Iterate Moving Average (IMA) viewpoint of SHB, we propose and analyze three novel step-size selections: MomSPSmax, MomDecSPS, and MomAdaSPS. For MomSPSmax, we provide convergence guarantees for SHB to a neighborhood of the solution for convex and smooth problems (without assuming interpolation). If interpolation is also satisfied, then using MomSPSmax, SHB converges to the true solution at a fast rate matching the deterministic HB. The other two variants, MomDecSPS and MomAdaSPS, are the first adaptive step-sizes for SHB that guarantee convergence to the exact minimizer without prior knowledge of the problem parameters and without assuming interpolation. The convergence analysis of SHB is tight and obtains the convergence guarantees of SGD with stochastic Polyak step-sizes as a special case. We supplement our analysis with experiments that validate the theory and demonstrate the effectiveness and robustness of the new algorithms.
@@ 1 Introduction
where each $f_{i}:\mathbb{R}^{d}\to\mathbb{R}$ is convex, smooth, and lower bounded by $\ell_{i}^{*}$. Let $X^{*}$ be the set of minimizers of (1). We assume that $X^{*}\neq\emptyset$ and we fix $x^{*}\in X^{*}$. This problem is the cornerstone of machine learning tasks (Hastie et al., 2009), where $x$ corresponds to the model parameters, $f_{i}(x)$ represents the loss on the training point $i$, and the aim is to minimize the average loss $f(x)$ across training points.
When $n$ is large, stochastic gradient methods are the preferred methods for solving (1) mainly because of their cheap per iteration cost. One of the most popular stochastic algorithms for solving such large-scale machine learning optimization problems is stochastic gradient descent (SGD) with momentum (Sutskever et al., 2013), a.k.a. stochastic heavy ball method (SHB) given by:
@@
where $S_{t}\subseteq[n]$ a random subset of data-points (mini-batch) with cardinality $B$ sampled independently at each iteration $t$, and $\nabla f_{S_{t}}(x^{t})=\frac{1}{B}\sum_{i\in S_{t}}\nabla f_{i}(x^{t})$ is the mini-batch gradient. Here $\gamma_{t}>0$ is the step-size/learning rate at iteration $t$ while $\beta_{t}\geq 0$ represents the momentum parameter. When the momentum parameter $\beta_{t}=0,\forall t\geq 0$, then the update rule SHB is equivalent to the well-studied mini-batch SGD, $x^{t+1}=x^{t}-\gamma_{t}\nabla f_{S_{t}}(x^{t})$ (Robbins and Monro, 1951), which has efficiently analyzed under different properties of problem (1) and different step-size selections $\gamma_{t}$ (Nemirovski and Yudin, 1983; Nemirovski et al., 2009; Hardt et al., 2016; Needell et al., 2016; Nguyen et al., 2018; Gower et al., 2019, 2021). Additionally, when the cardinality of the random subset $S_{t}$ is $B=n$, then the update rule of SHB is equivalent to the deterministic heavy ball method (HB) proposed by Polyak (1964), as a way to improve the convergence behavior of deterministic Gradient Descent (GD).
There is a rich literature on the convergence of SHB and HB in different scenarios. In Polyak (1964), it was proved that for a specific choice of the step-size $\gamma$ and the momentum parameter $\beta$, the HB method enjoys an accelerated linear convergence when minimizing strongly convex quadratic functions while more recently, Ghadimi et al. (2015) proved a global sublinear convergence guarantee for HB for convex and smooth functions. In the stochastic setting, several works focus on convergence guarantees of SHB under constant step-size and momentum parameters (Ma and Yarats, 2019; Kidambi et al., 2018; Yan et al., 2018; Gitman et al., 2019; Liu et al., 2020). However, in practical scenarios, these choices suffer from a prohibitively expensive and time-consuming hyper-parameter tuning process. This has motivated a large body of research on the development of adaptive SHB - a method that adapts their parameters using information collected during the iterative process. Such analysis is challenging, and the current adaptive versions of SHB either focus on the full batch setting (deterministic) (Barré et al., 2020; Saab Jr et al., 2022), or assume that an interpolation condition is satisfied (Wang et al., 2023; Schaipp et al., 2023b).
@@ 1.1 Main Contributions
$\diamond$ Efficient Polyak Step-sizes via IMA viewpoint.
@@
We explain and illustrate by experiment (see Figure 1) why naively using SPSmax of Loizou et al. (2021) as a step-size $\gamma_{t}$ in the update rule of SHB is not robust, leading to divergence even in simple problems. To resolve this issue, we provide an alternative way of selecting Polyak-type step-sizes for SHB via the Iterate Moving Average viewpoint from Sebbouh et al. (2021). Through our approach, we propose three novel adaptive step-size selections, namely MomSPSmax, MomDecSPS, MomAdaSPS. Each of the proposed step-sizes depends on the choice of the momentum parameter $\beta$ adding further stability to SHB and comes with specific benefits over their constant step-size counterparts or other adaptive variants of SHB.
$\diamond$ MomSPSmax: Convergence of SHB in non-interpolated setting.
@@
Our first step-size selection of SHB is MomSPSmax, which has a similar structure to SPSmax of Loizou et al. (2021) but includes also $(1-\beta)$ in its expression. For this choice, we provide convergence guarantees for SHB to a neighborhood of the solution for convex and smooth problems. Our analysis provides the first convergence guarantees of adaptive SHB using Polyak-type step-size. Previous works (Wang et al., 2023; Schaipp et al., 2023b) on Polyak step-size with momentum have guaranteed only under the interpolation setting. In addition, as a corollary of our main theoretical results, we show that MomSPSmax under the interpolation setting and in deterministic scenarios (full batch) converges to the true solution at a fast rate matching the deterministic HB.
$\diamond$ Convergence of SHB to exact solution via MomDecSPS and MomAdaSPS.
@@
Inspired by two recent Polyak-type step-size selections for SGD, the DecSPS in Orvieto et al. (2022) and AdaSPS in Jiang and Stich (2023), we propose two new ways for tuning the step-size for SHB. These are MomDecSPS and MomAdaSPS. Our analysis provides the first $O(1/\sqrt{T})$ convergence guarantees to the exact solution in the non-interpolated regime for a Polyak-type adaptive variant of SHB. Our proposed update rules converge for any choice of momentum parameter $\beta\in[0,1)$, which makes them particularly useful in practical scenarios.
$\diamond$ Tight Convergence Guarantees.
@@
All of our convergence guarantees are true generalizations of the theoretical analysis of SGD using SPSmax, DecSPS, and AdaSPS. That is, if $\beta=0$ (no momentum) in the update rule of SHB, our theorems obtain as a special case the best-known convergence rates of Polyak-type step-size for SGD, highlighting the tightness of our analysis. See also Table 1 for a summary of our main complexity results and a comparison with closely related works.
$\diamond$ Further Convergence Results.
@@
As a byproduct of our theoretical analysis, we provide two interesting corollaries: a novel analysis of constant step-size SHB (as a corollary of our Theorem on MomSPSmax) and the first robust convergence of SHB via our theorem on MomAdaSPS. For a constant step-size, the Corollary 3.4 of our Theorem 3.1 allows larger step-sizes than the analysis of SHB in Liu et al. (2020) and provides convergence without assuming the restrictive bounded variance condition (there exist $q>0$ such that $\operatorname{\mathbb{E}}\|\nabla f_{i}(x)-\nabla f(x)\|^{2}<q$). In addition, via Theorem 3.6, we provide the first robust convergence of adaptive SHB that guarantees convergence to the exact solution and automatically adapts to whether our problem is interpolated or not. That is, if interpolation is assumed, then the rate of SHB with MomAdaSPS is the same as the rate of SHB with MomSPSmax (or constant step-size), and if no interpolation is assumed, then it matches the rate of SHB with MomDecSPS. The analysis achieves the best-known rates in both settings.
$\diamond$ Numerical Evaluation.
@@
In Section 4, we verify our theoretical results via numerical experiments on various problems, demonstrating the effectiveness and practicality of our approach. An open-source implementation of our method is available at https://github.com/dimitris-oik/MomSPS.
[…]
@@
ALR-SMAG (Wang et al., 2023) Interpolation $O(\frac{1}{T}+\sigma^{2})$
MomSPSmax (Thm 3.1) Restriction on $\beta$ $O(\frac{1}{T}+\sigma^{2})$
MomDecSPS (Thm 3.5) Bounded Iterates $O(\frac{1}{\sqrt{T}})$
MomAdaSPS (Thm 3.6) Bounded Iterates $O(\frac{1}{T}+\frac{\sigma}{\sqrt{T}})$
[…]
Table 1: Summary of the considered step-sizes and the corresponding theoretical results. All the rates are given for convex and smooth functions. The quantity of convergence in our rates is $\operatorname{\mathbb{E}}[f(\overline{x}^{T})-f(x^{*})]$, where $\overline{x}^{T}=\frac{1}{T}\sum_{t=0}^{T}x^{t}$. Here $\hat{\sigma}^{2}=\operatorname{\mathbb{E}}\|\nabla f_{i}(x^{*})\|^{2}<\infty$ and $\sigma^{2}=\operatorname{\mathbb{E}}[f_{i}(x^{*})-\ell_{i}^{*}]<\infty$. The “Exact Convergence” column refers to convergence to the exact solution $x^{*}$ with no interpolation assumption.
@@ 2.2 Naive SPS in SHB
All of the above variants of SPS were proposed and analyzed for SGD (SHB with no momentum). With the increased popularity of momentum in machine learning, one can naturally ask the following question:
Is it possible to combine SPS with momentum?
This question was partially answered in Wang et al. (2023) when a momentum variant of SGD with Polyak-type step-size was proposed to solve the easier-to-analyze interpolated problems. In this scenario, the method behaves like its deterministic counterpart. However, several stochastic optimization problems do not satisfy the interpolation condition. In our work, we focus on this more general scenario.
@@
(a) $\beta=0.0$
(b) $\beta=0.2$
(c) $\beta=0.4$
(d) $\beta=0.5$
Figure 1: Comparison of MomSPSmax versus Naive SPSmax for various momentum coefficients $\beta$ on a logistic regression problem.
@@ 2.3 Iterative Moving Average: Balance between SPS and Momentum
Iterate Moving Average (IMA).
Sebbouh et al. (2021) provide the IMA as an alternative way of expressing the update rule of SHB and explain how the IMA formulation is crucial in comparing SHB and SGD as it allows to establish connections between the step-sizes of the two methods. The Iterate Moving Average method is given by the following update rule:
$z^{t+1}=z^{t}-\eta_{t}\nabla f_{S_{t}}(x^{t}),\quad x^{t+1}=\frac{\lambda_{t+1}}{\lambda_{t+1}+1}x^{t}+\frac{1}{\lambda_{t+1}+1}z^{t+1},$ (IMA)
@@
Three New Adaptive Step-sizes for SHB.
Using the IMA viewpoint and Proposition 2.1 and Corollary 2.2 let us present three Polyak-type step-size for SHB.
@@
where $\gamma_{-1}:=\gamma_{b}>0$ is a step-size bound and $c_{t}=c\sqrt{t+1}$ for $t\geq 0$ with $c_{-1}:=c_{0}=c>0$ be a constant to regulate the step-size. In this scenario we note that $\gamma_{t}$ is indeed decreasing since $\gamma_{t}\leq\frac{\gamma_{t-1}\sqrt{t}}{\sqrt{t+1}}\leq\gamma_{t-1}$. This step-size is adaptive and, in its definition, does not require knowledge of any function properties (e.g., smoothness/strong convexity constants). In Section 3 we prove that SHB with MomDecSPS and constant momentum $\beta\in[0,1)$ converges to exact solution with a rate $O(1/\sqrt{T})$.
@@
Let us close this section by mentioning two remarks that related to the above three step-size selections:
@@
The “correcting factor” $1-\beta$ is outside the minimum in the MomSPSmax while for the decreasing variants MomDecSPS and MomAdaSPS, it only appears in the first term of the minimum. This follows from Proposition 2.1 and the derivation of Corollary 2.2 and is explained in detail in Section C.3. In Section E.3, via experiments, we also test other variants of the above step-sizes with the correcting factor outside of the minimum. We observe that the other choices do not have good practical performance.

### v2

@@ Abstract
Stochastic gradient descent with momentum, also known as Stochastic Heavy Ball method (SHB), is one of the most popular algorithms for solving large-scale stochastic optimization problems in various machine learning tasks. In practical scenarios, tuning the step-size and momentum parameters of the method is a prohibitively expensive and time-consuming process. In this work, inspired by the recent advantages of stochastic Polyak step-size in the performance of stochastic gradient descent (SGD), we propose and explore new Polyak-type variants suitable for the update rule of the SHB method. In particular, using the Iterate Moving Average (IMA) viewpoint of SHB, we propose and analyze three novel step-size selections: MomSPSmax, MomDecSPS, and MomAdaSPS. For MomSPSmax, we provide convergence guarantees for SHB to a neighborhood of the solution for convex and smooth problems (without assuming interpolation). If interpolation is also satisfied, then using MomSPSmax, SHB converges to the true solution at a fast rate matching the deterministic HB. The other two variants, MomDecSPS and MomAdaSPS, are the first adaptive step-size for SHB that guarantee convergence to the exact minimizer - without a priori knowledge of the problem parameters and without assuming interpolation. Our convergence analysis of SHB is tight and obtains the convergence guarantees of stochastic Polyak step-size for SGD as a special case. We supplement our analysis with experiments validating our theory and demonstrating the effectiveness and robustness of our algorithms.
@@ 1 Introduction
where each $f_{i}:\mathbb{R}^{d}\to\mathbb{R}$ is convex, smooth, and lower bounded by $\ell_{i}^{*}$. Let $X^{*}$ be the set of minimizers of (1). We assume that $X^{*}\neq\emptyset$ and we fix $x^{*}\in X^{*}$. This problem is the cornerstone of machine learning tasks, (Hastie et al., 2009), where $x$ corresponds to the model parameters, $f_{i}(x)$ represents the loss on the training point $i$, and the aim is to minimize the average loss $f(x)$ across training points.
When $n$ is large, stochastic gradient methods are the preferred methods for solving (1) mainly because of their cheap per iteration cost. One of the most popular stochastic algorithms for solving such large-scale machine learning optimization problems is stochastic gradient descent (SGD) with momentum, (Sutskever et al., 2013), a.k.a. stochastic heavy ball method (SHB) given by:
@@
where $S_{t}\subseteq[n]$ a random subset of data-points (mini-batch) with cardinality $B$ sampled independently at each iteration $t$, and $\nabla f_{S_{t}}(x^{t})=\frac{1}{B}\sum_{i\in S_{t}}\nabla f_{i}(x^{t})$ is the mini-batch gradient. Here $\gamma_{t}>0$ is the step-size/learning rate at iteration $t$ while $\beta_{t}\geq 0$ represents the momentum parameter. When the momentum parameter $\beta_{t}=0,\forall t\geq 0$, then the update rule SHB is equivalent to the well-studied mini-batch SGD, $x^{t+1}=x^{t}-\gamma_{t}\nabla f_{S_{t}}(x^{t})$, (Robbins and Monro, 1951), which has been efficiently analyzed under different properties of problem (1) and different step-size selections $\gamma_{t}$ (Nemirovski and Yudin, 1983; Nemirovski et al., 2009; Hardt et al., 2016; Needell et al., 2016; Nguyen et al., 2018; Gower et al., 2019; 2021). Additionally, when the cardinality of the random subset $S_{t}$ is $B=n$, then the update rule of SHB is equivalent to the deterministic heavy ball method (HB) proposed by Polyak (1964), as a way to improve the convergence behavior of deterministic Gradient Descent (GD).
There is a rich literature on the convergence of SHB and HB in different scenarios. In Polyak (1964), it was proved that for a specific choice of the step-size $\gamma$ and the momentum parameter $\beta$, the HB method enjoys an accelerated linear convergence when minimizing strongly convex quadratic functions while more recently, Ghadimi et al. (2015) proved a global sublinear convergence guarantee for HB for convex and smooth functions. In the stochastic setting, several works focus on convergence guarantees of SHB under constant step-size and momentum parameters (Ma and Yarats, 2019; Kidambi et al., 2018; Yan et al., 2018; Gitman et al., 2019; Liu et al., 2020). However, in practical scenarios, these choices suffer from a prohibitively expensive and time-consuming hyper-parameter tuning process. This has motivated a large body of research on the development of adaptive SHB - a method that adapts their parameters using information collected during the iterative process. Such analysis is challenging, and the current adaptive versions of SHB either focus on the full batch setting (deterministic) (Barré et al., 2020; Saab Jr et al., 2022; Wang et al., 2023), or assume that an interpolation condition is satisfied (Schaipp et al., 2023b) or focus on moving averaged gradient (a different form of momentum) (Wang et al., 2023).
@@ 1.1 Main Contributions
$\diamond$ Efficient Polyak Step-sizes via IMA viewpoint. We explain and illustrate by experiment (see Figure 1) why naively using SPSmax of Loizou et al. (2021) as a step-size $\gamma_{t}$ in the update rule of SHB is not robust, leading to divergence even in simple problems. To resolve this issue, we provide an alternative way of selecting Polyak-type step-sizes for SHB via the Iterate Moving Average viewpoint from Sebbouh et al. (2021). Through our approach, we propose three novel adaptive step-size selections, namely MomSPSmax, MomDecSPS, MomAdaSPS. Each of the proposed step-sizes depends on the choice of the momentum parameter $\beta$ adding further stability to SHB and comes with specific benefits over their constant step-size counterparts or other adaptive variants of SHB.
@@
$\diamond$ MomSPSmax: Convergence of SHB in non-interpolated setting. Our first step-size selection of SHB is MomSPSmax, which has a similar structure to SPSmax of Loizou et al. (2021) but includes also $(1-\beta)$ in its expression. For this choice, we provide convergence guarantees for SHB to a neighborhood of the solution for convex and smooth problems. Our analysis provides the first convergence guarantees of adaptive SHB using Polyak-type step-size. Previous works on Polyak step-size with momentum in the stochastic setting have guarantees either only under the interpolation setting (Schaipp et al., 2023b) or for a moving averaged gradient momentum (Wang et al., 2023). In addition, as a corollary of our main theoretical results, we show that MomSPSmax under the interpolation setting and in deterministic scenarios (full batch) converges to the true solution at a fast rate matching the deterministic HB.
@@
$\diamond$ Convergence of SHB to exact solution via MomDecSPS and MomAdaSPS. Inspired by two recent Polyak-type step-size selections for SGD, the DecSPS in Orvieto et al. (2022) and AdaSPS in Jiang and Stich (2023), we propose two new ways for tuning the step-size for SHB. These are MomDecSPS and MomAdaSPS. Our analysis provides the first $O(1/\sqrt{T})$ convergence guarantees to the exact solution in the non-interpolated regime for a Polyak-type adaptive variant of SHB. Our proposed update rules converge for any choice of momentum parameter $\beta\in[0,1)$, which makes them particularly useful in practical scenarios.
@@
$\diamond$ Tight Convergence Guarantees. All of our convergence guarantees are true generalizations of the theoretical analysis of SGD using SPSmax, DecSPS, and AdaSPS. That is, if $\beta=0$ (no momentum) in the update rule of SHB, our theorems obtain as a special case the best-known convergence rates of Polyak-type step-size for SGD, highlighting the tightness of our analysis. See also Table 1 for a summary of our main complexity results and a comparison with closely related works.
@@
$\diamond$ Further Convergence Results. As a byproduct of our theoretical analysis, we provide two interesting corollaries: a novel analysis of constant step-size SHB (as a corollary of our Theorem on MomSPSmax) and the first robust convergence of SHB via our theorem on MomAdaSPS. For a constant step-size, the Corollary 3.5 of our Theorem 3.2 allows larger step-sizes than the analysis of SHB in Liu et al. (2020) and provides convergence without assuming the restrictive bounded variance condition (there exist $q>0$ such that $\operatorname{\mathbb{E}}\|\nabla f_{i}(x)-\nabla f(x)\|^{2}<q$). In addition, via Theorem 3.7 we provide the first robust convergence of adaptive SHB that guarantees convergence to the exact solution and automatically adapts to whether our problem is interpolated or not. That is, if interpolation is assumed, then the rate of SHB with MomAdaSPS is the same as the rate of SHB with MomSPSmax (or constant step-size), and if no interpolation is assumed, then it matches the rate of SHB with MomDecSPS. The analysis achieves the best-known rates in both settings.
@@
$\diamond$ Numerical Evaluation. In Section 4, we verify our theoretical results via numerical experiments on various problems, demonstrating the effectiveness and practicality of our approach. An open-source implementation of our method is available at https://github.com/dimitris-oik/MomSPS.
Step-size Assumptions / Setting Adaptive Exact Convergence Rate
@@
ALR-SMAG (Wang et al., 2023) Moving Averaged Gradient $O(\frac{1}{T}+\sigma^{2})$
MomSPSmax (Thm 3.2) Restriction on $\beta$ $O(\frac{1}{T}+\sigma^{2})$
MomDecSPS (Thm 3.6) Bounded Iterates $O(\frac{1}{\sqrt{T}})$
MomAdaSPS (Thm 3.7) Bounded Iterates $O(\frac{1}{T}+\frac{\sigma}{\sqrt{T}})$
Table 1: Summary of the considered step-sizes and the corresponding theoretical results in the stochastic setting. All the rates are given for convex and smooth functions. The quantity of convergence in our rates is $\operatorname{\mathbb{E}}[f(\overline{x}^{T})-f(x^{*})]$, where $\overline{x}^{T}=\frac{1}{T}\sum_{t=0}^{T}x^{t}$. Here $\hat{\sigma}^{2}=\operatorname{\mathbb{E}}\|\nabla f_{i}(x^{*})\|^{2}<\infty$ and $\sigma^{2}=\operatorname{\mathbb{E}}[f_{i}(x^{*})-\ell_{i}^{*}]<\infty$. The “Exact Convergence” column refers to convergence to the exact solution $x^{*}$ with no interpolation assumption.
@@ 2.2 Naive SPS in SHB
All of the above variants of SPS were proposed and analyzed for SGD (SHB with no momentum). With the increased popularity of momentum in machine learning, one can naturally ask the following question: Is it possible to combine SPS with momentum? This question was partially answered in Wang et al. (2023) when they proposed a Polyak-type step-size for a moving averaged gradient (MAG) momentum but not for the SHB. Their MAG framework is not equivalent to SHB when their stepsizes are adaptive and while their proposed stepsize reduces to the original SPSmax stepsize when $\beta=0$, their guarantees do not reduce to the original SPSmax guarantees for convex functions.
Figure 1: Comparison of MomSPSmax versus SPSmax with naive momentum for different momentum parameters $\beta$ on a logistic regression problem. Left: $\beta=0.2$, Right: $\beta=0.5$
@@ 2.3 Iterative Moving Average: Balance between SPS and Momentum
Iterate Moving Average (IMA). Sebbouh et al. (2021) provide the IMA as an alternative way of expressing the update rule of SHB and explain how the IMA formulation is crucial in comparing SHB and SGD as it allows to establish connections between the step-sizes of the two methods. The Iterate Moving Average method is given by the following update rule:
$z^{t+1}=z^{t}-\eta_{t}\nabla f_{S_{t}}(x^{t}),\quad x^{t+1}=\frac{\lambda_{t+1}}{\lambda_{t+1}+1}x^{t}+\frac{1}{\lambda_{t+1}+1}z^{t+1},$
@@
Three New Adaptive Step-sizes for SHB. Using the IMA viewpoint and Proposition 2.1 and Corollary 2.2 let us present three Polyak-type step-size for SHB.
@@
where $\gamma_{-1}:=\gamma_{b}>0$ is a step-size bound and $c_{t}=c\sqrt{t+1}$ for $t\geq 0$ with $c_{-1}:=c_{0}=c>0$ be a constant to regulate the step-size. Note that $\gamma_{t}$ is indeed decreasing. This step-size is adaptive and, in its definition, does not require knowledge of any function properties (e.g., smoothness/strong convexity constants). In Section 3 we prove that SHB with MomDecSPS and constant momentum $\beta\in[0,1)$ converges to exact solution with a rate $O(1/\sqrt{T})$.
@@
Let us close this section by mentioning two remarks related to the above three step-size selections:
@@
The “correcting factor” $1-\beta$ is outside the minimum in the MomSPSmax while for the decreasing variants MomDecSPS and MomAdaSPS, it only appears in the first term of the minimum. This follows from Proposition 2.1 and the derivation of Corollary 2.2 and is explained in detail in Section C.3. In Section F.3, via experiments, we also test other variants of the above step-sizes with the correcting factor outside of the minimum. We observe that the other choices do not have good practical performance.